# DexterCap: An Affordable and Automated System for Capturing Dexterous Hand-Object Manipulation

### v1

@@ Abstract.
Hand-Object Interaction, Motion Capture, Dataset, Animation
copyright: noneccs: Computing methodologies Motion capture
Figure 1. DexterCap captures dexterous manipulation of a Rubik’s Cube. Top: raw multi-camera footage showing character-coded marker patches. Bottom: reconstructed 3D hand and object motion. Dense marker coverage enables accurate reconstruction of subtle, natural in-hand manipulation motions.
1. Introduction
@@
Despite its importance, the precise capture and realistic modeling of fine-grained hand skills remain a significant challenge. Although human-object interaction (HOI) datasets and models are becoming increasingly available, most existing work focuses on basic actions such as grasping and moving, or interactions with simple articulated objects (Taheri et al., 2020; Fan et al., 2023; Kim et al., 2025). There is a substantial gap in resources for understanding and synthesizing the nuanced dexterity required for a wide range of hand skills.
@@
Our system draws inspiration from the vision-based approach introduced by (Chen et al., 2021a) for capturing human body deformation during full-body motion. They demonstrated that deep learning-based computer vision techniques, combined with specially designed visual patterns, enables the automated identification and localization of a dense set of markers. Building on this idea for dexterous hand-object manipulation, we present a practical, low-cost methodology that operates with affordable hardware and minimal labeling, applies to arbitrary rigid objects, and is readily extendable to articulated ones. By leveraging a large number of tagged markers, a system can enhance robustness against occlusion and significantly reduce the need for tedious manual cleanup.
@@
We address the first challenge by introducing a three-stage marker detection and identification model that cascades the detection of markers, edges, and tags, enabling robust and efficient marker recognition. To tackle the second challenge, we employ the morphable hand model MANO (Romero et al., 2017) and develop a robust calibration and solving procedure to ensure accurate reconstruction of both hand and object poses.
@@
2. Related Work
2.1. Motion Capture for Human-Object Interaction
@@
Marker-based optical motion capture is often considered the high-precision “gold standard”, with systems like Qualisys and Vicon widely used. These systems use multiple cameras to track markers, enabling sub-millimeter-accurate 3D reconstruction from high-frame-rate positional data (Lu et al., 2025a; Kim et al., 2025; Han et al., 2018). Despite their accuracy, these systems are costly and prone to occlusion issues (e.g., marker swapping), which often necessitates extensive manual post-processing to ensure the quality of captured motion data. Novel solutions include T-pose initialization (Meyer et al., 2014) and learning-based correspondence methods that use synthetic data and neural networks to improve marker solving robustness (Holden, 2018; Han et al., 2018; Chen et al., 2021b; Ghorbani and Black, 2021; Pan et al., 2023). Other approaches leverage monocular video cues (Milef et al., 2024) or employ unique markers (Chen et al., 2021a). However, when extending such systems to more interaction scenarios, particularly those involving multi-jointed objects like a Rubik’s cube, a tedious manual process is still required.
@@
Sensor-based methods, using Inertial Measurement Units (IMUs) or joint force contact sensors, are less prone to visual occlusion and more cost-effective. To further improve the usability, recent research has explored reducing the number of IMUs in TransPose (Yi et al., 2021) and PIP (Yi et al., 2022), and enhancing accuracy by mitigating issues such as local inertial forces (Yi et al., 2024). However, capturing fine-grained hand motions with IMU gloves still remains challenging in terms of finger accuracy. Moreover, in interaction-rich scenarios, the introduction of objects typically requires additional markers and alignment  (Kim et al., 2025).
@@
Vision-based markerless methods estimate body or hand pose directly from images or videos, eliminating the need for markers or specialized suits and thus offering greater convenience and lower cost (El Moubtahij et al., 2022). These approaches leverage various imaging modalities, including single/multiple RGB cameras (Pavlakos et al., 2024; Potamias et al., 2025), depth cameras (Hu et al., 2022), infrared cameras (Baker et al., 2025)), with multi-view systems alleviating occlusion and enhancing 3D accuracy. Despite these advances, vision-based markerless methods still face persistent challenges, including severe occlusions, lighting variations, motion blur, and maintaining temporal consistency (El Moubtahij et al., 2022). Furthermore, although there are recent works that incorporate reinforcement learning to enforce realistic hand-object interactions (Hu et al., 2024), ensuring the physical plausibility of the estimated poses still remains difficult.
2.2. Hand-Object Interaction Datasets
Object manipulation datasets are crucial for advancing robotics and human-computer interaction, as they address challenges like complex hand movements, varied object shapes, and dynamic contacts, enabling effective model training and evaluation.
@@ 2.2. Hand-Object Interaction Datasets
Existing human-object interaction datasets primarily focus on capturing coordinated whole-body interactions with objects, but often lack precise details regarding hand-object contact  (Kratzer et al., 2020; Bhatnagar et al., 2022; Li et al., 2023; Yang et al., 2024; Zhang et al., 2023; Araújo et al., 2023). The integration of optical motion capture systems has enabled the collection of more detailed grasping motions  (Taheri et al., 2020; Tendulkar et al., 2023; Fu et al., 2025). Recent work expands these datasets to include a wider variety of bimanual functional manipulations  (Lv et al., 2024) and interactions with articulated objects  (Fan et al., 2023; Zhan et al., 2024; Lu et al., 2025b). Instead of focusing on whole-body interaction with broad hand motion, our dataset specifically targets fine-grained in-hand manipulation.
@@
To capture hand-object interactions, many datasets rely on RGB-D reconstruction, which, while generalizable, often falls short in recording detailed hand movements  (Hasson et al., 2019; Hampali et al., 2020; Chao et al., 2021; Kwon et al., 2021; Liu et al., 2022; Yang et al., 2022). Datasets such as ContactPose  (Brahmbhatt et al., 2020) emphasize the collection of affordance  (Corona et al., 2020) and contact information, which can improve grasp accuracy but still lack dynamic hand motion details. While ManipNet  (Zhang et al., 2021) includes precise in-hand manipulations, such as twisting and rotating  (Moon et al., 2020), its object coverage remains limited to rigid items. Thanks to our system design and specialized marker patches, our dataset captures accurate in-hand contact and subtle hand motions. Additionally, we include complex object interactions, such as those involving a Rubik’s cube, in our dataset.
3. Dexterous Hand Motion Capture System
@@
Our system, DexterCap, is a multi-camera, vision-based motion capture and reconstruction framework designed to achieve high-precision tracking of hand and object motions in dexterous in-hand interactions. It offers a relatively low-cost solution with an easily deployable and automated hardware-software setup. The system consists of a set of synchronized industrial cameras positioned around the capture space from multiple directions. We attach a set of specially designed visual markers to the subject’s hand and the object. Hand motions are captured simultaneously by all cameras. These visual markers are then detected in the recorded videos with minimal manual labeling, aided by learning-based marker detection models. Finally, hand and finger motions are estimated using the morphable hand model MANO (Romero et al., 2017), while object poses are recovered using object-specific solvers. In this section, we detail the system design and associated algorithms.
3.1. Hardware Setup and Marker Design
3.1.1. Cameras
@@
Our current DexterCap setup uses a set of synchronized industrial cameras that record grayscale video at a resolution of $2048\times{}2448$ and 20 FPS. The cameras are mounted on a $2\times{}1\times{}2$ m cage and are arranged to minimize occlusions and ensure robust 3D hand tracking. All cameras are calibrated using the standard procedure with the OpenCV library (Bradski, 2000). A short exposure time of 1ms is used to prevent motion blur, and the aperture is narrowed as much as possible to increase depth of field. The system is installed in a bright room with natural lighting, supplemented by a small set of LED lights that provide additional illumination when needed.
3.1.2. Markers
@@
Figure 2. Our marker system: (a) Markers are attached to rigid hand regions (phalanges, dorsum) for accurate tracking. (b) Top: Rigid object (green cube) with visual marker patches. Bottom left: A visual marker patch printed on a transfer sticker. Bottom right: Visual markers transferred onto a medical tape for secure attachment.
Inspired by (Chen et al., 2021a), we use a customized marker system with high-contrast checkerboard patterns for robust tracking and identification. As illustrated in Figure 2, each white square contains a unique two-character ID in black, selected from 26 uppercase letters and 10 digits, excluding visually similar characters, yielding 324 unique tags. We place an underscore below the left character to help determine orientation.
@@
Figure 3. Image processing pipeline for marker detection. (a) Raw input image with character-coded checkerboard markers. (b) CornerNet detection results. (c) EdgeNet edge classification results. (d) BlockNet block recognition with character identifiers.
(Chen et al., 2021a) created a special suit with visual patterns printed directly onto the fabric. However, we found that the equivalent idea using a glove does not work well for hand motion capture. A glove can stretch and wrinkle under different hand poses and may slide along the skin, which introduces errors in pose estimation. Instead, we attach marker patches directly to each relatively rigid region of the hand, including the finger knuckles, the dorsum, and the palm, for a total of 19 patches per hand. These patches are transfer printed onto medical adhesive tape for reliable and comfortable attachment. This dense marker layout provides over 500 detectable corners, ensuring accurate motion reconstruction even under severe occlusion. Markers are also placed on object surfaces to support robust pose estimation for both rigid and articulated objects such as a Rubik’s Cube.
3.2. Image Processing and Marker Extraction
Once the raw video data is acquired, each frame is processed through our image processing pipeline to detect and identify visual markers. As illustrated in Figure 3, the pipeline first detects candidate corners, then identifies edges between them to assemble blocks, and finally recognizes the tag of each block, which encode the corner IDs. Commonly used marker systems, such as ArUco (Garrido-Jurado et al., 2014) from OpenCV, often perform poorly in this setting due to severe deformation of the markers. To address this, we train dedicated models for marker detection.
3.2.1. Corner Detection
Precise localization of the corners in the checkerboard-style marker pattern is critical to our system. To achieve this, we adopt a deep learning approach based on the U-Net (Ronneberger et al., 2015) architecture, which we refer to as CornerNet in our work. CornerNet takes a $64\times 64$ grayscale image as input and produces a heatmap of the same resolution, where each corner is represented by a Gaussian peak. The raw input images are divided into overlapping patches of $64\times 64$ pixel for efficient processing. For details of CornerNet, please refer to Appendix C. To handle scale variation, we also downsample the images by half and process them in the same manner, which provides an expanded field of view and improves robustness to perspective changes. Compared with direct coordinate regression (Chen et al., 2021a), this heatmap-based formulation is more reliable at both patch and image levels (see Sec. 5.2). It also allows each patch to contain multiple corners and avoids ambiguities when merging overlapping predictions.
@@ 3.2.1. Corner Detection
3.2.2. Edge and Block Detection
With the candidate corners detected, the next step is to determine how they assemble into checkerboard blocks. (Chen et al., 2021a) proposed validating all candidate quads generated based on geometric criteria. However, we find these criteria difficult to configure, and they often produce a large number of candidate quads, resulting in high computational cost during block validation. To address this, we adopt a two-stage block detection strategy. First, we check whether each pair of candidate corners is directly connected in the template pattern, and then assemble blocks based on these connections.
Specifically, we employ a neural network, EdgeNet, based on ResNet (He et al., 2016) with a fully connected classification head that acts as a binary classifier to determine whether a valid edge exists between a pair of candidate corners. For each pair of candidate corners within a predefined Euclidean distance threshold, we extract a rectangular image patch centered on the segment connecting the two corners and aligned with its direction. The distance threshold is determined from the statistics of edge lengths in the manually annotated training set and is chosen to achieve 95% recall. This choice balances accuracy and computational efficiency by excluding implausibly long connections while preserving the majority of valid edges. The extracted patch is resized to $64\times 64$ and used as input to EdgeNet. Once all connections are validated, we perform a graph search to identify corners and edges that form convex quadrilateral blocks with acceptable aspect ratios, avoiding shapes that are excessively thin or distorted.
@@ 3.2.2. Edge and Block Detection
This edge-first assembly prunes obviously invalid candidates early and avoids redundant computation on shared edges, shrinking candidates by an order of magnitude and improving precision/recall over (Chen et al., 2021a) (see Sec. 5.2).
3.2.3. Block Identification
The final stage of the image processing pipeline is to validate and identify the detected blocks from the previous steps. Following (Chen et al., 2021a), we employ a neural network, BlockNet, for this task. Our BlockNet is a ResNet-based architecture with three output heads, each responsible for predicting one character of the tag and the orientation of the block. We use one-hot encoding for both the tag letters and the orientation, which allows BlockNet to be effectively trained using the standard cross-entropy loss. At inference time, blocks that are determined to be invalid, along with their corresponding corners, are discarded.
3.2.4. Post-Processing
@@
3.4. Hand Motion Reconstruction
With the 3D markers reconstructed from the previous step, we proceed to solve for hand and object motions. We use the parametric MANO model (Romero et al., 2017) as the hand template. In this section, we describe how we adapt the model, perform calibration, and solve for hand motion.
3.4.1. Hand Model Parameterization

### v2

# DexterCap: Affordable and Automated Capture of Complex Hand-Object Interactions

@@ Abstract
{CCSXML}
<ccs2012> <concept> <concept_id>10010147.10010371.10010352.10010238</concept_id> <concept_desc>Computing methodologies Motion capture</concept_desc> <concept_significance>500</concept_significance> </concept> </ccs2012>
[…]
[500]Computing methodologies Motion capture
[…]
volume: 45issue: 2
1 Introduction
@@
Despite its importance, the precise capture and realistic modeling of fine-grained hand skills remain a significant challenge. Although human-object interaction (HOI) datasets and models are becoming increasingly available, most existing work focuses on basic actions such as grasping and moving, or interactions with simple articulated objects [taheri2020grab, fan2023arctic, kim2024parahome]. There is a substantial gap in resources for understanding and synthesizing the nuanced dexterity required for a wide range of hand skills.
@@
Our system draws inspiration from the vision-based approach introduced by [chen2021capturing] for capturing human body deformation during full-body motion. They demonstrated that deep learning-based computer vision techniques, combined with specially designed visual patterns, enables the automated identification and localization of a dense set of markers. Building on this idea for dexterous hand-object manipulation, we present a practical, low-cost methodology that operates with affordable hardware and minimal labeling, applies to arbitrary rigid objects, and is readily extendable to articulated ones. By leveraging a large number of tagged markers, a system can enhance robustness against occlusion and significantly reduce the need for tedious manual cleanup.
@@
We address the first challenge by introducing a three-stage marker detection and identification model that cascades the detection of markers, edges, and tags, enabling robust and efficient marker recognition. To tackle the second challenge, we employ the morphable hand model MANO [romero2017embodied] and develop a robust calibration and solving procedure to ensure accurate reconstruction of both hand and object poses.
@@
2 Related Work
2.1 Motion Capture for Human-Object Interaction
@@
Marker-based optical motion capture is often considered the high-precision “gold standard”, with systems like Qualisys and Vicon widely used. These systems use multiple cameras to track markers, enabling sub-millimeter-accurate 3D reconstruction from high-frame-rate positional data [lu2025humoto, kim2024parahome, han2018online]. Despite their accuracy, these systems are costly and prone to occlusion issues (e.g., marker swapping), which often necessitates extensive manual post-processing to ensure the quality of captured motion data. Novel solutions include T-pose initialization [meyer2014online] and learning-based correspondence methods that use synthetic data and neural networks to improve marker solving robustness [holden2018robust, han2018online, chen2021mocap, ghorbani2021soma, pan2023locality]. Other approaches leverage monocular video cues [milef2024towards] or employ unique markers [chen2021capturing]. However, when extending such systems to more interaction scenarios, particularly those involving multi-jointed objects like a Rubik’s cube, a tedious manual process is still required.
@@
Sensor-based methods, using Inertial Measurement Units (IMUs) or joint force contact sensors, are less prone to visual occlusion and more cost-effective. To further improve the usability, recent research has explored reducing the number of IMUs in TransPose [yi2021transpose] and PIP [yi2022physical], and enhancing accuracy by mitigating issues such as local inertial forces [yi2024physical]. However, capturing fine-grained hand motions with IMU gloves still remains challenging in terms of finger accuracy. Moreover, in interaction-rich scenarios, the introduction of objects typically requires additional markers and alignment  [kim2024parahome].
@@
Vision-based markerless methods estimate body or hand pose directly from images or videos, eliminating the need for markers or specialized suits and thus offering greater convenience and lower cost [elmoubtahij2022structured]. These approaches leverage various imaging modalities, including single/multiple RGB cameras [pavlakos2024reconstructing, potamias2024wilor], depth cameras [hu2022physical], infrared cameras [baker2025optimisation], with multi-view systems alleviating occlusion and enhancing 3D accuracy. Despite these advances, vision-based markerless methods still face persistent challenges, including severe occlusions, lighting variations, motion blur, and maintaining temporal consistency [elmoubtahij2022structured]. Furthermore, although there are recent works that incorporate reinforcement learning to enforce realistic hand-object interactions [haoyu2024hoic], ensuring the physical plausibility of the estimated poses still remains difficult.
2.2 Hand-Object Interaction Datasets
@@
Existing human-object interaction datasets primarily focus on capturing coordinated whole-body interactions with objects, but often lack precise details regarding hand-object contact  [MOGAZE, BEHAVE, OMOMO, SEMANTIC-HOI, HODome, CIRCLE]. The integration of optical motion capture systems has enabled the collection of more detailed grasping motions  [taheri2020grab, FLEX, fu2025gigahands]. Recent work expands these datasets to include a wider variety of bimanual functional manipulations  [HIMO] and interactions with articulated objects  [fan2023arctic, OAKINK2, CHOICE]. Instead of focusing on whole-body interaction with broad hand motion, our dataset specifically targets fine-grained in-hand manipulation.
@@
To capture hand-object interactions, many datasets rely on RGB-D reconstruction, which, while generalizable, often falls short in recording detailed hand movements  [OBMAN, hampali2020honnotate, DexYCB, H2O, HOI4D, OakInk]. Datasets such as ContactPose  [CONTACTPOSE] emphasize the collection of affordance  [YCB-AFFORDANCE] and contact information, which can improve grasp accuracy but still lack dynamic hand motion details. While ManipNet  [MANIPNET] includes precise in-hand manipulations, such as twisting and rotating  [INTERHAND26M], its object coverage remains limited to rigid items. Thanks to our system design and specialized marker patches, our dataset captures accurate in-hand contact and subtle hand motions. Additionally, we include complex object interactions, such as those involving a Rubik’s cube, in our dataset.
3 Dexterous Hand Motion Capture System
@@
Our system, DexterCap, is a multi-camera, vision-based motion capture and reconstruction framework designed to achieve high-precision tracking of hand and object motions in dexterous in-hand interactions. It offers a relatively low-cost solution with an easily deployable and automated hardware-software setup. The system consists of a set of synchronized industrial cameras positioned around the capture space from multiple directions. We attach a set of specially designed visual markers to the subject’s hand and the object. Hand motions are captured simultaneously by all cameras. These visual markers are then detected in the recorded videos with minimal manual labeling, aided by learning-based marker detection models. Finally, hand and finger motions are estimated using the morphable hand model MANO [romero2017embodied], while object poses are recovered using object-specific solvers. In this section, we detail the system design and associated algorithms.
3.1 Hardware Setup and Marker Design
3.1.1 Cameras
@@
Our current DexterCap setup uses a set of synchronized industrial cameras that record grayscale video at a resolution of $2048\times{}2448$ and 20 FPS. The cameras are mounted on a $2\times{}1\times{}2$ m cage and are arranged to minimize occlusions and ensure robust 3D hand tracking. All cameras are calibrated using the standard procedure with the OpenCV library [opencv_library]. A short exposure time of 1ms is used to prevent motion blur, and the aperture is narrowed as much as possible to increase depth of field. The system is installed in a bright room with natural lighting, supplemented by a small set of LED lights that provide additional illumination when needed.
3.1.2 Markers
@@
Figure 1: Our marker system: (a) Markers are attached to rigid hand regions (phalanges, dorsum) for accurate tracking. (b) Top: Rigid object (green cube) with visual marker patches. Bottom left: A visual marker patch printed on a transfer sticker. Bottom right: Visual markers transferred onto a medical tape for secure attachment.
Inspired by [chen2021capturing], we use a customized marker system with high-contrast checkerboard patterns for robust tracking and identification. As illustrated in Figure 1, each white square contains a unique two-character ID in black, selected from 26 uppercase letters and 10 digits, excluding visually similar characters, yielding 324 unique tags. We place an underscore below the left character to help determine orientation.
@@
Figure 2: Image processing pipeline for marker detection. (a) Raw input image with character-coded checkerboard markers. (b) CornerNet detection results. (c) EdgeNet edge classification results. (d) BlockNet block recognition with character identifiers.
[chen2021capturing] created a special suit with visual patterns printed directly onto the fabric. However, we found that the equivalent idea using a glove does not work well for hand motion capture. A glove can stretch and wrinkle under different hand poses and may slide along the skin, which introduces errors in pose estimation. Instead, we attach marker patches directly to each relatively rigid region of the hand, including the finger knuckles, the dorsum, and the palm, for a total of 19 patches per hand. These patches are transfer printed onto medical adhesive tape for reliable and comfortable attachment. This dense marker layout provides over 500 detectable corners, ensuring accurate motion reconstruction even under severe occlusion. Markers are also placed on object surfaces to support robust pose estimation for both rigid and articulated objects such as a Rubik’s Cube.
3.2 Image Processing and Marker Extraction
Once the raw video data is acquired, each frame is processed through our image processing pipeline to detect and identify visual markers. As illustrated in Figure 2, the pipeline first detects candidate corners, then identifies edges between them to assemble blocks, and finally recognizes the tag of each block, which encode the corner IDs. Commonly used marker systems, such as ArUco [ArUco] from OpenCV, often perform poorly in this setting due to severe deformation of the markers. To address this, we train dedicated models for marker detection.
3.2.1 Corner Detection
Precise localization of the corners in the checkerboard-style marker pattern is critical to our system. To achieve this, we adopt a deep learning approach based on the U-Net [ronneberger2015unet] architecture, which we refer to as CornerNet in our work. CornerNet takes a $64\times 64$ grayscale image as input and produces a heatmap of the same resolution, where each corner is represented by a Gaussian peak. The raw input images are divided into overlapping patches of $64\times 64$ pixel for efficient processing. For details of CornerNet, please refer to Appendix C. To handle scale variation, we also downsample the images by half and process them in the same manner, which provides an expanded field of view and improves robustness to perspective changes. Compared with direct coordinate regression [chen2021capturing], this heatmap-based formulation is more reliable at both patch and image levels (see Sec. 5.2). It also allows each patch to contain multiple corners and avoids ambiguities when merging overlapping predictions.
@@ 3.2.1 Corner Detection
3.2.2 Edge and Block Detection
With the candidate corners detected, the next step is to determine how they assemble into checkerboard blocks. [chen2021capturing] proposed validating all candidate quads generated based on geometric criteria. However, we find these criteria difficult to configure, and they often produce a large number of candidate quads, resulting in high computational cost during block validation. To address this, we adopt a two-stage block detection strategy. First, we check whether each pair of candidate corners is directly connected in the template pattern, and then assemble blocks based on these connections.
Specifically, we employ a neural network, EdgeNet, based on ResNet [he2015resnet] with a fully connected classification head that acts as a binary classifier to determine whether a valid edge exists between a pair of candidate corners. For each pair of candidate corners within a predefined Euclidean distance threshold, we extract a rectangular image patch centered on the segment connecting the two corners and aligned with its direction. The distance threshold is determined from the statistics of edge lengths in the manually annotated training set and is chosen to achieve 95% recall. This choice balances accuracy and computational efficiency by excluding implausibly long connections while preserving the majority of valid edges. The extracted patch is resized to $64\times 64$ and used as input to EdgeNet. Once all connections are validated, we perform a graph search to identify corners and edges that form convex quadrilateral blocks with acceptable aspect ratios, avoiding shapes that are excessively thin or distorted.
@@ 3.2.2 Edge and Block Detection
This edge-first assembly prunes obviously invalid candidates early and avoids redundant computation on shared edges, shrinking candidates by an order of magnitude and improving precision/recall over [chen2021capturing] (see Sec. 5.2).
3.2.3 Block Identification
The final stage of the image processing pipeline is to validate and identify the detected blocks from the previous steps. Following [chen2021capturing], we employ a neural network, BlockNet, for this task. Our BlockNet is a ResNet-based architecture with three output heads, each responsible for predicting one character of the tag and the orientation of the block. We use one-hot encoding for both the tag letters and the orientation, which allows BlockNet to be effectively trained using the standard cross-entropy loss. At inference time, blocks that are determined to be invalid, along with their corresponding corners, are discarded.
3.2.4 Post-Processing
@@
3.4 Hand Motion Reconstruction
With the 3D markers reconstructed from the previous step, we proceed to solve for hand and object motions. We use the parametric MANO model [romero2017embodied] as the hand template. In this section, we describe how we adapt the model, perform calibration, and solve for hand motion.
3.4.1 Hand Model Parameterization